OpenStudy (anonymous):

if the sides of a square are lengthened by 8cm. the area becomes144cm^2. find the length of a side of the original square

OpenStudy (anonymous):

(x+8)^2 = 144 (x+8) = 12 x = 12-8 = 4

OpenStudy (anonymous):

the area is one of the sides, squared (because it's a square). so you know that length of one side of that square is 12, or the square root of 144. subtracting 8 from this leaves 4.